# Approximating the set of Nash equilibria for convex games

### v2

@@ 4.1 Computing a polyhedral approximation of the upper image
Algorithm 1 of [11] allows for the computation of a polyhedral $\varepsilon$-approximation (10) to problem (3) for a direction $c_{i}\in\operatorname{int}\mathbb{R}^{m_{i}}_{+}$. Since, by Theorem 3.2, we have to use $c_{i}=\bar{c}_{i}:=(0,....,0,1)^{\top}\notin\operatorname{int}\mathbb{R}^{m_{i}}_{+}$ instead, we cannot apply [11, Algorithm 1] directly. A careful inspection of [11, Algorithm 1] reveals that the assumption $c_{i}\in\operatorname{int}\mathbb{R}^{m_{i}}_{+}$ is only used in [11, Proposition 4.4] to show that the so-called Pascoletti-Serafini scalarization of problem (3), that is,
@@
for $v\in\mathbb{R}^{m_{i}}$, is feasible and strong duality is satisfied between $(P_{2}(v))$ and its Lagrange dual problem. This in turn is used to prove the correctness of [11, Algorithm 1] in [11, Theorem 4.9]. For the direction $\bar{c}_{i}=(0,....,0,1)^{\top}$, feasibility and strong duality might fail for the points $v$ considered in the course of the algorithm. We will now introduce a modification to [11, Algorithm 1] that allows for the specific problem of interest, problem (3), to restore feasibility and strong duality for $(P_{2}(v))$ and its dual. This modification takes advantage of the fact that the zero entries of $\bar{c}_{i}$ correspond to linear components in the objective of problem (3).
Let us now explain the details of this modification. Algorithm 1 of [11] starts with an initialization phase where an initial outer approximation $\mathcal{P}_{0}$ of $\mathcal{P}_{i}$ is computed via solving weighted sum scalarizations. The solutions of the corresponding weighted sum scalarizations are added to an initial set $\bar{\mathcal{X}}$. This is done in line 2 and 3 of [11, Algorithm 1] and remains the same in the modified version. It is shown in [11] that the initial outer approximation $\mathcal{P}_{0}$ is pointed.
However, before entering the iteration phase in line 4 of [11, Algorithm 1], we will add the following two lines to the algorithm, where we denote by $a_{i}:=\frac{{m_{i}}-1}{2}$:
@@
Recall that, by Assumption 4.2, $\mathbb{X}=\operatorname{conv}\{\bar{x}^{1},...,\bar{x}^{k}\}$ for $k\in\mathbb{N}$. This completes the now modified initialization phase of the algorithm. The rest of the algorithm (i.e. the iteration phase of [11, Algorithm 1]) remains unchanged.
@@
Let us now comment on these two modifications. Line (11) makes additional cuts to the outer approximation $\mathcal{P}_{0}$. The first cut with the linear subspace $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}=-y_{a_{i}+1:m_{i}-1}\}$ is using the particular structure of problem (3), utilizing that the first $a_{i}$ components of the objective function are the negative of the next $a_{i}$ components. Thus, the upper image has to lie in this linear subspace and this fact can be used for the initial outer approximation $\mathcal{P}_{0}$ already. The second cut with the polyhedron $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}\in\operatorname{conv}\{\bar{x}^{1}_{-i},...,\bar{x}^{k}_{-i}\}\}$ is using that, by Assumption 4.2, $\mathbb{X}=\operatorname{conv}\{\bar{x}^{1},...,\bar{x}^{k}\}$ for $k\in\mathbb{N}$ and the fact that the first $a_{i}$ components of the objective function of problem (3) are $x_{-i}$. Since both, the linear space $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}=-y_{a_{i}+1:m_{i}-1}\}$ and the polyhedron $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}\in\operatorname{conv}\{\bar{x}^{1}_{-i},...,\bar{x}^{k}_{-i}\}\}$, are supersets of the image set $\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}$ of (3), the updated set $\mathcal{P}_{0}$ remains still an outer approximation of this image set. (Note that one could add +$\mathbb{R}^{m_{i}}_{+}$ to $\mathcal{P}_{0}$ obtained in (11) in order to obtain an $\mathcal{P}_{0}$ that would still be an outer approximation of $\mathcal{P}_{i}$, but since it is enough for it to be an outer approximation of the image set, the addition of +$\mathbb{R}^{m_{i}}_{+}$ is not needed for the algorithm to work correctly.) Further, the updated set $\mathcal{P}_{0}$ is still a pointed polyhedron. Line (12) updates the set $\bar{\mathcal{X}}$ by adding the vertices of the feasible set $\mathbb{X}$. This does not affect [11, Algorithm 1] or its correctness and only changes the set $\bar{\mathcal{X}}$ outputted by the algorithm. It is however crucial for step 3 of Algorithm 4.7 introduced below, see also Section 4.3, as it has an important consequence as detailed in Remark 4.5.
@@
In the second phase of [11, Algorithm 1] the outer approximation $\mathcal{P}_{0}$ is iteratively updated until after termination of the algorithm, (10) is satisfied. We prove in the following that the iteration phase of [11, Algorithm 1] works also correctly for problem (3) with respect to the boundary direction $\bar{c}_{i}$, if the modification of the initialization phase, i.e. the one with the two additional lines (11) and (12), is used.
@@
The idea of the iteration phase of [11, Algorithm 1] is that in each iteration it is checked if the distance of the vertices of the current outer approximation $\mathcal{P}_{0}$ to the upper image $\mathcal{P}_{i}$ is less or equal than $\varepsilon$. If for each vertex the distance is less or equal than $\varepsilon$ then the algorithm stops. If a vertex of $\mathcal{P}_{0}$ is found to have a distance larger than $\varepsilon$, then the set $\mathcal{P}_{0}$ is updated. To check for each vertex $v\in\mathcal{P}_{0}$ the distance to $\mathcal{P}_{i}$, the Pascoletti-Serafini scalarization $(P_{2}(v))$ is considered. Since $v$ is always chosen among the vertices of $\mathcal{P}_{0}$, we can, using (11), rewrite $(P_{2}(v))$ as
@@
Consider $v\in\mathcal{P}_{0}$. By (11) we know that $v_{1:a_{i}}\in\operatorname{conv}\{\bar{x}^{1}_{-i},...,\bar{x}^{k}_{-i}\}$. Thus, there are $\lambda^{1},...,\lambda^{k}\geq 0$ with $\sum_{j=1}^{k}\lambda^{j}=1$ such that $v_{1:a_{i}}=\sum_{j=1}^{k}\lambda^{j}\bar{x}^{j}_{-i}$. Set $\bar{x}=\sum_{j=1}^{k}\lambda^{j}\bar{x}^{j}\in\mathbb{X}$ and choose $\bar{z}\in\mathbb{R}$ such that $f_{i}(\bar{x})-\bar{z}-v_{m_{i}}\leq 0$. Then $(\bar{x},\bar{z})$ is feasible for $(P_{2}(v))$. Since the feasible set of $(P_{2}(v))$ is compact (by the same arguments as in the proof of [11, Proposition 4.4]), an optimal solution exists. In order to guarantee strong duality we need to find a feasible element of $(P_{2}(v))$ that satisfies the weak Slater condition, i.e. for all nonaffine constraints the inequality constraint needs to be strict. However, since the only nonaffine constraint of $(P_{2}(v))$ is $f_{i}({x})-{z}-v_{m_{i}}\leq 0$ we can always find $z^{*}\in\mathbb{R}$ such that $f_{i}(\bar{x})-z^{*}-v_{m_{i}}<0$. Then $(\bar{x},z^{*})$ satisfies the weak Slater condition which implies strong duality. ∎
This result guarantees (by replacing in the proof of [11, Theorem 4.9], statement [11, Proposition 4.4] by Lemma 4.4) that the iteration phase works correctly under the changes done in the initialization phase. Thus, the modification of [11, Algorithm 1] proposed here outputs, at termination, a finite set $\bar{\mathcal{X}}$ such that $\mathcal{P}^{In}_{i}:=\operatorname{conv}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\bar{\mathcal{X}}\}+\mathbb{R}^{m_{i}}_{+}$ satisfying (10).
@@
Denote the elements of the finite set $\bar{\mathcal{X}}$ by $\{{x}^{1},...,{x}^{s}\}$ for $s\in\mathbb{N}$ and consider $\mathcal{P}^{In}_{i}:=\operatorname{conv}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\bar{\mathcal{X}}\}+\mathbb{R}^{m_{i}}_{+}$. Since for any vertex $\bar{x}$ of $\mathbb{X}$, we have, due to (12), $(\bar{x}_{-i},-\bar{x}_{-i},f_{i}(\bar{x}))\in\mathcal{P}^{In}_{i}$, we conclude that for any $x\in\mathbb{X}$ it holds $x_{-i}\in\operatorname{conv}\{{x}^{1}_{-i},...,{x}^{s}_{-i}\}$. As such, for any Pareto optimal element $x^{*}$ of (3) we can express $x^{*}_{-i}$ via convex combinations of ${x}^{1}_{-i},...,{x}^{s}_{-i}$. This is necessary to prove the correctness of Algorithm 4.7 proposed in Subsection 4.4.
@@ 4.2 Computing the maximal efficient faces
4.3 Approximate the set of Pareto optimal points
The next step is to approximate the set of Pareto optimal points of problem (3). In detail, we want to approximate the set of preimages whose images lie below the maximal efficient faces $F^{j}$, $j=1,...,k_{i}$, computed in Section 4.2, yielding a set of $\epsilon$-Pareto optimal points $X_{i}$ satisfying (9).
The aim is to compute
@@ 4.3 Approximate the set of Pareto optimal points
for each maximal efficient face $F^{j}$, where $j=1,...,k_{i}$. Note that this is a bounded convex projection problem as considered in [8, 9] as it is of the form: compute $\text{compute }\{x\in\mathbb{X}\;|\;\exists y\in F^{j}:(x,y)\in S\}$ for $S=\{(x,y)\in\mathbb{X}\times F^{j}\;|\;(x_{-i},-x_{-i},f_{i}(x))\leq_{\mathbb{R}^{m_{i}}_{+}}y\}$. The set $S$ is bounded due to Assumption 4.2 and Remark 4.6. As the set $S$ is convex (and not polyhedral in general), the sets $\bar{X}^{j}_{i}$ cannot be computed exactly, but can only be approximated by the methods in e.g. [19, 12, 8] or [9, Algorithm 5.1]. Note that [19, 12, 8] solve an associated convex multi-objective problem in dimension $d+1$ for $d=\sum_{j=1}^{N}n_{j}$, whereas [9, Algorithm 5.1] works in dimension $d$. For a fixed error level $\varepsilon_{2}>0$ one obtains a finite set $\hat{X}^{j}_{i}\subseteq\mathbb{X}$ satisfying
@@
see e.g. [8, Theorem 3.11] and [9, Theorem 5.4], where $B_{\varepsilon_{2}}$ is a closed $\varepsilon_{2}$-ball around the origin (in the $L_{1}$ norm). In detail, [9, Algorithm 5.1] outputs a finite $\varepsilon_{2}$-solution $\bar{S}\subseteq S$ of the bounded convex projection problem (14) such that $\hat{X}^{j}_{i}:=\{x\in\mathbb{X}\;|\;(x,y)\in\bar{S}\}$ (i.e. the collection of the $x$ components of the finitely many vectors $(x,y)\in\bar{S}$) satisfies (15).
@@
is the desired set of $\epsilon$-Pareto optimal points of problem (3) satisfying (9) for an adapted error level $\epsilon$, that involves the two error levels $\varepsilon_{1},\varepsilon_{2}>0$ and the Lipschitz constant of the function $f_{i}$. Hence the set
$X=\bigcap_{i=1}^{N}X_{i}$
will be the desired set yielding (8):

### v4

@@ 4.1 Computing a polyhedral approximation of the upper image
Algorithm 1 and 2 of [14] allow for the computation of a polyhedral $\varepsilon$-approximation (10) to problem (3) for a direction $c_{i}\in\operatorname{int}\mathbb{R}^{m_{i}}_{+}$. Since, by Theorem 3.2, we have to use $c_{i}=\bar{c}_{i}:=(0,....,0,1)^{\top}\notin\operatorname{int}\mathbb{R}^{m_{i}}_{+}$ instead, we cannot apply [14, Algorithm 1 or 2] directly. Let us consider Algorithm 1 of [14] first. Algorithm 2 of [14] is considered in Remark 4.6 below. A careful inspection of [14, Algorithm 1] reveals that the assumption $c_{i}\in\operatorname{int}\mathbb{R}^{m_{i}}_{+}$ is only used in [14, Proposition 4.4] to show that the so-called Pascoletti-Serafini scalarization of problem (3) for $v\in\mathbb{R}^{m_{i}}$, that is,
@@
is feasible and strong duality is satisfied between $(P_{2}(v))$ and its Lagrange dual problem. This in turn is used to prove the correctness of [14, Algorithm 1] in [14, Theorem 4.9]. For the direction $\bar{c}_{i}=(0,....,0,1)^{\top}$, feasibility and strong duality might fail for the points $v$ considered in the course of the algorithm. We will now introduce a modification to [14, Algorithm 1] that allows for the specific problem of interest, problem (3), to restore feasibility and strong duality for $(P_{2}(v))$ and its dual. This modification takes advantage of the fact that the zero entries of $\bar{c}_{i}$ correspond to linear components in the objective of problem (3).
Let us now explain the details of this modification. Algorithm 1 of [14] starts with an initialization phase where an initial outer approximation $\mathcal{P}_{0}$ of $\mathcal{P}_{i}$ is computed via solving weighted sum scalarizations. The solutions of the corresponding weighted sum scalarizations are added to an initial set $\bar{\mathcal{X}}$. This is done in line 2 and 3 of [14, Algorithm 1] and remains the same in the modified version. It is shown in [14] that the initial outer approximation $\mathcal{P}_{0}$ is pointed.
However, before entering the iteration phase in line 4 of [14, Algorithm 1], we will add the following two lines to the algorithm, where we denote by $a_{i}:=\frac{{m_{i}}-1}{2}$:
@@
Recall that $\mathbb{X}=\operatorname{conv}\{\bar{x}^{1},...,\bar{x}^{k}\}$ for $k\in\mathbb{N}$ (Assumption 4.2). This completes the now modified initialization phase. The rest of the algorithm (i.e. the iteration phase of [14, Algorithm 1]) remains unchanged.
@@
Let us now comment on these two modifications. Line (11) makes additional cuts to the outer approximation $\mathcal{P}_{0}$. The first cut with the linear subspace $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}=-y_{a_{i}+1:m_{i}-1}\}$ is using the particular structure of problem (3), utilizing that the first $a_{i}$ components of the objective function are the negative of the next $a_{i}$ components. Thus, the upper image has to lie in this linear subspace and this fact can be used for the initial outer approximation $\mathcal{P}_{0}$ already. The second cut with the polyhedron $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}\in\operatorname{conv}\{\bar{x}^{1}_{-i},...,\bar{x}^{k}_{-i}\}\}$ is using that, by Assumption 4.2, $\mathbb{X}=\operatorname{conv}\{\bar{x}^{1},...,\bar{x}^{k}\}$ for $k\in\mathbb{N}$ and the fact that the first $a_{i}$ components of the objective function of problem (3) are $x_{-i}$. Since both, the linear space $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}=-y_{a_{i}+1:m_{i}-1}\}$ and the polyhedron $\{y\in\mathbb{R}^{m_{i}}\;|\;y_{1:a_{i}}\in\operatorname{conv}\{\bar{x}^{1}_{-i},...,\bar{x}^{k}_{-i}\}\}$, are supersets of the image set $\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\mathbb{X}\}$ of (3), the updated set $\mathcal{P}_{0}$ remains still an outer approximation of this image set. (Note that one could add +$\mathbb{R}^{m_{i}}_{+}$ to $\mathcal{P}_{0}$ obtained in (11) in order to obtain an $\mathcal{P}_{0}$ that would still be an outer approximation of $\mathcal{P}_{i}$, but since it is enough for it to be an outer approximation of the image set, the addition of +$\mathbb{R}^{m_{i}}_{+}$ is not needed for the algorithm to work correctly.) Further, the updated set $\mathcal{P}_{0}$ is still a pointed polyhedron. Line (12) updates the set $\bar{\mathcal{X}}$ by adding the vertices of the feasible set $\mathbb{X}$. This does not affect [14, Algorithm 1] or its correctness and only changes the set $\bar{\mathcal{X}}$ outputted by the algorithm. It is however crucial for step 3 of Algorithm 4.8 introduced below, see also Section 4.3, as it has an important consequence as detailed in Remark 4.5.
@@
In the second phase of [14, Algorithm 1] the outer approximation $\mathcal{P}_{0}$ is iteratively updated until after termination of the algorithm, (10) is satisfied. We prove in the following that the iteration phase of [14, Algorithm 1] works also correctly for problem (3) with respect to the boundary direction $\bar{c}_{i}$, if the modification of the initialization phase, i.e. the one with the two additional lines (11) and (12), is used.
@@
The idea of the iteration phase of [14, Algorithm 1] is that in each iteration it is checked if the distance of the vertices of the current outer approximation $\mathcal{P}_{0}$ to the upper image $\mathcal{P}_{i}$ is less or equal than $\varepsilon$. If for each vertex the distance is less or equal than $\varepsilon$ then the algorithm stops. If a vertex of $\mathcal{P}_{0}$ is found to have a distance larger than $\varepsilon$, then the set $\mathcal{P}_{0}$ is updated. To check for each vertex $v\in\mathcal{P}_{0}$ the distance to $\mathcal{P}_{i}$, the Pascoletti-Serafini scalarization $(P_{2}(v))$ is considered. Since $v$ is always chosen among the vertices of $\mathcal{P}_{0}$, we can, using (11), rewrite $(P_{2}(v))$ as
@@
Consider $v\in\mathcal{P}_{0}$. By (11) we know that $v_{1:a_{i}}\in\operatorname{conv}\{\bar{x}^{1}_{-i},...,\bar{x}^{k}_{-i}\}$. Thus, there are $\lambda^{1},...,\lambda^{k}\geq 0$ with $\sum_{j=1}^{k}\lambda^{j}=1$ such that $v_{1:a_{i}}=\sum_{j=1}^{k}\lambda^{j}\bar{x}^{j}_{-i}$. Set $\bar{x}=\sum_{j=1}^{k}\lambda^{j}\bar{x}^{j}\in\mathbb{X}$ and choose $\bar{z}\in\mathbb{R}$ such that $f_{i}(\bar{x})-\bar{z}-v_{m_{i}}\leq 0$. Then $(\bar{x},\bar{z})$ is feasible for $(P_{2}(v))$. Since the feasible set of $(P_{2}(v))$ is compact (by the same arguments as in the proof of [14, Proposition 4.4]), an optimal solution exists. In order to guarantee strong duality we need to find a feasible element of $(P_{2}(v))$ that satisfies the weak Slater condition, i.e. for all nonaffine constraints the inequality constraint needs to be strict. However, since the only nonaffine constraint of $(P_{2}(v))$ is $f_{i}({x})-{z}-v_{m_{i}}\leq 0$ we can always find $z^{*}\in\mathbb{R}$ such that $f_{i}(\bar{x})-z^{*}-v_{m_{i}}<0$. Then $(\bar{x},z^{*})$ satisfies the weak Slater condition which implies strong duality. ∎
This result guarantees (by replacing in the proof of [14, Theorem 4.9], statement [14, Proposition 4.4] by Lemma 4.4) that the iteration phase works correctly under the changes done in the initialization phase. Thus, the modification of [14, Algorithm 1] proposed here outputs, at termination, a finite set $\bar{\mathcal{X}}$ such that $\mathcal{P}^{In}_{i}:=\operatorname{conv}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\bar{\mathcal{X}}\}+\mathbb{R}^{m_{i}}_{+}$ satisfies (10).
@@
Denote the elements of the finite set $\bar{\mathcal{X}}$ produced at termination by the modification of [14, Algorithm 1] by $\{{x}^{1},...,{x}^{s}\}$ for $s\in\mathbb{N}$ and consider $\mathcal{P}^{In}_{i}:=\operatorname{conv}\{(x_{-i},-x_{-i},f_{i}(x))\;|\;x\in\bar{\mathcal{X}}\}+\mathbb{R}^{m_{i}}_{+}$. Since for any vertex $\bar{x}$ of $\mathbb{X}$, we have, due to (12), $(\bar{x}_{-i},-\bar{x}_{-i},f_{i}(\bar{x}))\in\mathcal{P}^{In}_{i}$, we conclude that for any $x\in\mathbb{X}$ it holds $x_{-i}\in\operatorname{conv}\{{x}^{1}_{-i},...,{x}^{s}_{-i}\}$. As such, for any Pareto optimal element $x^{*}$ of (3) we can express $x^{*}_{-i}$ via convex combinations of ${x}^{1}_{-i},...,{x}^{s}_{-i}$. This is necessary to prove the correctness of Algorithm 4.8 proposed in Subsection 4.4.
Remark 4.6.
Let us now consider the dual algorithm, Algorithm 2 of [14]. The assumption $c_{i}\in\operatorname{int}\mathbb{R}^{m_{i}}_{+}$ is only used in [14, Proposition 3.10], which is needed to show the correctness of this algorithm. One can prove that for the considered convex multi-objective optimization problem (3), [14, Proposition 3.10] holds also for direction $c_{i}=\bar{c}_{i}:=(0,....,0,1)^{\top}\notin\operatorname{int}\mathbb{R}^{m_{i}}_{+}$. Hence, Algorithm 2 of [14] can be applied directly to problem (3). However, the addition of line (12) is also needed here as it is crucial for step 3 of Algorithm 4.8 introduced below.
@@ 4.2 Computing the maximal efficient faces
4.3 Approximating the set of Pareto optimal points
The next step is to approximate the set of Pareto optimal points of problem (3). In detail, we want to approximate the set of preimages whose images lie below the maximal efficient faces $F^{j}$, $j=1,...,k_{i}$, computed in Section 4.2, yielding a set of $\epsilon$-Pareto optimal points $X_{i}$ satisfying (9). The aim is to compute
@@ 4.3 Approximating the set of Pareto optimal points
for each maximal efficient face $F^{j}$, where $j=1,...,k_{i}$. Note that this is a bounded convex projection problem as considered in [10, 11] as it is of the form: compute $\text{compute }\{x\in\mathbb{X}\;|\;\exists y\in F^{j}:(x,y)\in S\}$ for $S=\{(x,y)\in\mathbb{X}\times F^{j}\;|\;(x_{-i},-x_{-i},f_{i}(x))\leq_{\mathbb{R}^{m_{i}}_{+}}y\}$. The set $S$ is bounded due to Assumption 4.2 and Remark 4.7. As the set $S$ is convex (and not polyhedral in general), the sets $\bar{X}^{j}_{i}$ cannot be computed exactly, but can only be approximated by the methods in e.g. [22, 15, 10] or [11, Algorithm 5.1]. Note that [22, 15, 10] solve an associated convex multi-objective problem in dimension $d+1$ for $d=\sum_{j=1}^{N}n_{j}$, whereas [11, Algorithm 5.1] works in dimension $d$. For a fixed error level $\varepsilon_{2}>0$ one obtains a finite set $\hat{X}^{j}_{i}\subseteq\mathbb{X}$ satisfying
@@
see e.g. [10, Theorem 3.11] and [11, Theorem 5.4], where $B_{\varepsilon_{2}}$ is a closed $\varepsilon_{2}$-ball around the origin (in the $L_{1}$ norm). In detail, [11, Algorithm 5.1] outputs a finite $\varepsilon_{2}$-solution $\bar{S}\subseteq S$ of the bounded convex projection problem (14) such that $\hat{X}^{j}_{i}:=\{x\in\mathbb{X}\;|\;(x,y)\in\bar{S}\}$ (i.e. the collection of the $x$ components of the finitely many vectors $(x,y)\in\bar{S}$) satisfies (15).
@@
is the desired subset of $\epsilon$-Pareto optimal points of problem (3) satisfying (9) for an adapted error level $\epsilon$, that involves the two error levels $\varepsilon_{1},\varepsilon_{2}>0$ and the Lipschitz constant of the function $f_{i}$. Hence the set $X=\bigcap_{i=1}^{N}X_{i}$ will be the desired set yielding (8):